OpenStudy (anonymous):

how do I find the answer to 40% =2 divided by B

myininaya (myininaya):

.4=2/B .4B=2 B=2/.4=20/4=5

OpenStudy (anonymous):

OKAY but what does B equal? I have to graph it. Needs to be a number ?? Maybe I don't get what you did.

myininaya (myininaya):

you don't see B=5?

OpenStudy (anonymous):

I have to do ten of these and sorry but I'm OLD and didn't quite follow HOW you got the answer and no I didn't see it sorry. I sort of see the steps of how to solve it.

myininaya (myininaya):

40%=.4 .4=2/B multiply both sides by B .4B=2 divide both sides by .4 B=2/.4 B=2 divided by .4 B=2 divided by 4/10 B=2 divided by 2/5 =2*5/2=5 because the twos cancel

OpenStudy (anonymous):

Thank you for being so explicit LOL I need it !!

myininaya (myininaya):

your welcome if you want you can give me a medal

OpenStudy (anonymous):

I'm sorry but I still don't get it'

myininaya (myininaya):

which part?

myininaya (myininaya):

40%=.4=4/10=2/5 maybe this is easier for you to understand so 2/5=2/B how can these fractions be equal what does B have to be in order for the equation to be true

OpenStudy (anonymous):

OK I tried one tell me if I got it right please 52%=A divided by 25 Is A 13 ?

myininaya (myininaya):

52%=.52=52/100=13/25 13/25=A/25 yes A would have to be 13 gj

OpenStudy (anonymous):

OK I think I can handle it .. Thanks =))